OpenStudy (anonymous):

All of the following are ways to write 25 percent of N EXCEPT A. 0.25 N B. 25N/100 C. (1\4)N D. 25 N

OpenStudy (anonymous):

a. 0.25N

OpenStudy (anonymous):

why ??

OpenStudy (anonymous):

how you get it ?!!

OpenStudy (anonymous):

25 percent of N = ( 25/100 ) N => never greater than N

OpenStudy (anonymous):

0.25 = 25/100 = 1/4

OpenStudy (anonymous):

so it will be A or C ??

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

okay but how you get it ??

OpenStudy (anonymous):

wait its D

OpenStudy (anonymous):

ok but how ?? :D

OpenStudy (anonymous):

got mixed yo, its D because 25 N 25 x N and well because thats not a percentage :) the otherrs are correct.

OpenStudy (anonymous):

you get it? :p

OpenStudy (anonymous):

hhhhhhhhhhhh no ^_^

OpenStudy (anonymous):

ok i'll break it down. A. is correct because if you divide 1/4 = 0.25 so choice A is correct. B. 25/100 = 0.25 so that is correct

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Letter C. (1/4) = 0.25 so that is correct choice. Lastly letter D, which is 25 n... you would need to know what N equals to. so that one is wrong . because it would be way higher than 25.

OpenStudy (anonymous):

you are looking for 0.25% nothing higher than that, so D would be the wrong one.

OpenStudy (anonymous):

ahaaa get it know ^_^

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

welcome :) good!